OpenStudy (princesssleelee):

The function for the cost of materials to make a shirt is f(x) = five sixths x + 5, where x is the number of shirts. The function for the selling price of those shirts is g(f(x)), where g(x) = 5x + 6. Find the selling price of 18 shirts. 106 120 185 196

OpenStudy (pawanyadav):

OK......what is the value of f(x) for x=18?

OpenStudy (4everaddicted2anime):

the answer is 106

OpenStudy (princesssleelee):

Im thinking the overall answer is 106

OpenStudy (princesssleelee):

Im not sure though

OpenStudy (pawanyadav):

How you get the answer? ... Show your work I'll check if there is any mistake.

OpenStudy (4everaddicted2anime):

5((5/6)*18+5)+6 5(20)+6 100+6 =106

OpenStudy (princesssleelee):

f(x)=56x+5 g(x)=5x+6 so at x = 18 I need to solve f(x) first when x = 18 so f(x) 56x+5 when x = 18 we have 56(18)+5 equaling 20 take that 20 and plug it into g(x) for every 'x' 5x+6 will become 5(20)+6 equaling 106 :)